The charges and coordinates of two charged particles held fixed in an xy plane are q1 = 3.49 μC, x1 = 3.80 cm, y1 = 0.165 cm and q2 = -5.05 μC, x2 = -2.87 cm, y2 = 2.08 cm. Find the (a)magnitude and (b) direction (with respect to +x-axis in the range (-180°;180°]) of the electrostatic force on particle 2 due to particle 1. At what (c) x and (d) y coordinates should a third particle of charge q3 = 3.77 μC be placed such that the net electrostatic force on particle 2 due to particles 1 and 3 is zero?Chapter 21, Problem 015 Your answer is partially correct. Try again. The charges and coordinates of two charged particles held fixed in an xy plane are 91-3.49 μC, X1-3.80 cm, y1-0.165 cm and q2--5.05 μC, X2 -2.87 cm, y2-2.08 cm. Find the (a) magnitude and (b) direction (with respect to +x-axis in the range (-180°;180°]) of the electrostatic force on particle 2 due to particle 1. At what (c) x and (d) y coordinates should a third particle of charge q3-3.77 μC be placed such that the net electrostatic force on particle 2 due to particles 1 and 3 is zero? (a) Number (b) Number (c) Number (d) N Click if you would like to Show Work for this question: UnitsN 47.4 Units T (degrees) 0821 Units m 1.231 Uni Open Show Work
Answer #1

r = sqrt ((-2.87 - 3.80)2 + (2.08 - 0.165)2)

r = sqrt ((-6.67)2 + (1.915)2)

r = 6.939 cm

or

r = 0.06939 m

Therefore

F = 9e9 * 3.49e-6 * 5.05e-6 / 0.069392

F = 32.943 N

--------------------------------------------------------------------------

Direction

theta = arctan (1.915 / -6.67)

theta = -16.019 degree

-------------------------------------------------------------

setting force between q2 and q3 equal to force between q2 and q1, we have

q3 / r12 = q1 / r22

Therefore,

r12 = sqrt (3.77*6.9392 / 3.49)

r1 = 7.2119 cm

x = -2.87 - 7.2119* cos16.019

x = -9.8018 cm or - 0.098018 m

y = 2.08 + 7.2119*sin 16.019

y = 4.07 cm or 0.0407 m